1. What happens to the temperature of a substance as it changes from a liquid to a gas?



  1. Calculate what mass of water can be changed to steam if 10.6kJ of heat energy is supplied to the water at 1000C. (Specific latent heat of vaporization of water = 2.26x106J/kg)






  1. Ammonia is vaporised in order to freeze an ice rink.

a)  Find out how much heat it would take to vaporise 1 g of ammonia.

    (Specific latent heat of vaporisation of ammonia = 1.34 x 10J/kg      Specific latent heat of fusion of ice = 3.34 x 10J/kg).


1
Expert's answer
2021-05-25T10:09:44-0400

The temperature of a substance does not change when a liquid becomes a gas.

The mass of water can be changed to steam is


"m=\frac QL=\frac{10600}{2.26·10^6}=0.00469\text{ kg}."

Find out how much heat it would take to vaporise 1 g of ammonia:


"Q=mL=0.001·1.34·10^6=1.34·10^3\text{ J}."